OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)^2 – 25? Identify the x–intercepts by graphing the quadratic equation.

OpenStudy (anonymous):

let x = 0 y-intercept = (0-3^2 - 25 = -16

OpenStudy (anonymous):

y=x^2+9-25 would come out to be equation simplified

OpenStudy (anonymous):

let y = 0 0 = (x -3)^ - 25 or 25 = (x-3)^2 take the square root of both sides 5 = x - 3 or -5 = x - 3 solve x = 8 or x = -2

OpenStudy (anonymous):

well thats thee quadratic equation

OpenStudy (anonymous):

and follow the link to find the graph http://www.wolframalpha.com/input/?i=solve+0%3D%28x+%E2%80%93+3%29^2+%E2%80%93+25